OpenStudy (anonymous):

evaluate w^3 when w=4

OpenStudy (anonymous):

64

OpenStudy (anonymous):

64

OpenStudy (angela210793):

4^3=4*4*4=16*4 =64

OpenStudy (anonymous):

4^3=64